OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10?

OpenStudy (mathstudent55):

Solve for y, so you'll have it in the y = mx + b form, and m is the slope.

OpenStudy (anonymous):

alternatively, know that the slope of \[ax+by=c\] is always \(-\frac{a}{b}\)

OpenStudy (anonymous):

You can easily find the slope by changing it from standard form (ax + by = c) to slope intercept form (y= mx + b) Just carry over the x term to the right side, and isolate the y 2x - 5y = -10 - 2x -2x ------------- -5y = -2x -10 ------------- -5 y = 2/5x - 10 This is now slope intercept form --> y= mx +b , where m is the slope m= 2/5

OpenStudy (anonymous):

thank you that helps a lot !!!!

OpenStudy (anonymous):

You're welcome